OpenStudy (anonymous):

Determine which system below will produce infinitely many solutions. 2x + 5y = 24 2x + 5y = 42 ------- 3x - 2y = 15 6x + 5y = 11 ------- 4x - 3y = 9 -8x + 6y = -18 ------- 5x - 3y = 16 -2x + 3y = -7

OpenStudy (anonymous):

@iGreen @sammixboo can one of yo help me by helping me start off solving the first set if equations? after that I will see about the others and chose my answers n stuff

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Well, for it to have infinitely many solutions, they have to be the same line.

OpenStudy (anonymous):

how do I know if they are the same line? is it if the equations are the same, but the answer is different? or what

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (igreen):

Well if we multiply the whole equation by 1 number, it will stay the same. Like say we have: 3x + 5y = 2 If we multiply this by 6 we get: 18x + 30y = 12 This will be the same line as 3x + 5y = 2.

OpenStudy (anonymous):

took me a while but A is parrelel cause they have the same expression but not the same answer

OpenStudy (anonymous):

ugh

OpenStudy (whpalmer4):

yes, the two lines given by \[2x+5y=24\]and\[2x+5y=42\]are parallel and so there are no solutions to that system of equations, as solutions occur only where all of the functions intersect.

OpenStudy (igreen):

Yep, you got it.

OpenStudy (igreen):

Now for B. Divide 6 / 3 to see if the 2nd equation is the first equation multiplied by a similar number.

OpenStudy (igreen):

I got 6 and 3 from 6x and 3x.

OpenStudy (anonymous):

ok i know that can you help be solve for B?

OpenStudy (whpalmer4):

it is maybe a bit clearer if you consider rearranging them into slope intercept form: \[2x+5y=24\]solve for \(y\)\[5y=24-2x\]\[y=-\frac{2}{5}x+\frac{24}5\] \[2x+5y=42\]becomes \[y=-\frac{2}{5}y+\frac{42}{5}\]by similar algebra Clearly, for any given value of \(x\), the second one gives you a value of \(y\) that is more positive.

OpenStudy (anonymous):

B is 3x - 2y = 15 6x + 5y = 11

OpenStudy (igreen):

OS is so slow

OpenStudy (anonymous):

i was hoping to use the elimination method, but after I add n stuff idk what to do next

OpenStudy (whpalmer4):

So, if the two lines have different slopes, there will be 1 solution, because they are not parallel and they are not identical, so they will intersect somewhere. That somewhere is the 1 solution. If the lines have identical slopes, then you need to check the y-intercepts. If the y-intercepts are identical, the two lines are identical and there are infinitely many solutions, as any point on the line works, and a line is made up of an infinite number of points. If the lines have identical slopes by differing y-intercepts, then they are parallel, and there are no solutions.

OpenStudy (anonymous):

see we have 3x - 2y = 15 6x + 5y = 11 add 3x + 6x to get 9x then add -2y + 5y to get 3y then add 15 + 11 to get 26 to we have 9x + 3y = 26

OpenStudy (igreen):

No..that's not how we do it..

OpenStudy (anonymous):

or do we multiply -2 by everything in the first equation to get -6x + 4y = -30 6x + 5y = 11

OpenStudy (anonymous):

then we add them together using elimination then get 0 + 9y = -19

OpenStudy (igreen):

That's to get the solution, yes.

OpenStudy (igreen):

@KaylaIsBae Try multiplying -2 to the first equation for Option C.

OpenStudy (whpalmer4):

another way to check is to try multiplying one equation through by a constant to see if you can make the coefficients match those of the other equation. For example: \[3x+2y = 6\]\[-6x-4y=-12\]If we multiply the first equation by \(-2\), that \(3x\) becomes \(-6x\), just like we have in the other equation. What happens to the rest of it? \[(-2)3x+(-2)2y=(-2)6\]\[-6x-4y=-12\] We've converted it to be an exact copy of the second equation. Therefore, these two lines are identical. Another example: \[3x+2y=6\]\[-6x-4y=2\]Multiply first equation by \(-2\) giving \[(-2)3x+(-2)2y=(-2)6\]\[-6x-4y=-12\]That has the same \(x\) and \(y\) coefficients as the second equation, but a different constant. Those lines are parallel. Remember that the (negative) ratio of the \(x&y\) coefficients forms the slope: \[Ax+By=C\]\[By=-Ax+C\]\[y=-(\frac{A}{B})x+\frac{C}{A}\] which when compared with slope intercept form \[y=mx+b\]shows the slope and y-intercept.

OpenStudy (anonymous):

i am only supposed to use substitution method or elimination method

OpenStudy (whpalmer4):

Yikes, all those posts interspersed with my stuff...none of that showed up while I was posting!

OpenStudy (anonymous):

seethat makes sense but i am still confused. i know what the answer is too, i wanna know how to use these

OpenStudy (anonymous):

methods

OpenStudy (anonymous):

for B imma try the substitution method tell me if i am wrong 3x - 2y = 15 6x + 5y = 11 taking the first question imma subtract 3x on both sides 3x = 15 + 2y now dividing 3 on both sides x = 15/3 + 2y/3 x = 5 + 2y/3

OpenStudy (anonymous):

doesn't 2y/3 = 2/3y

OpenStudy (anonymous):

uuuuuuuuuuuuuuuuuuuuuuuuuuuuuugh